OpenStudy (anonymous):

Give the domain and range of f(x)=(x-5)^3. What is the domain and range of f(x)? Will the answer be all real numbers or am I missing something?

OpenStudy (anonymous):

domain: x values of f(x) with sense range: f(x) values

OpenStudy (anonymous):

In this case: domain \[-\infty < x < \infty\] range \[-\infty < y < \infty\]

OpenStudy (anonymous):

Oh wow, thanks.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."